OpenStudy (anonymous):

How do I balance C2H16+O2=

OpenStudy (rane):

ist can u show me your products as well

OpenStudy (anonymous):

carbondioxide and water vapour (CO2 + H2O)

OpenStudy (rane):

write the full equation

OpenStudy (anonymous):

C2H16+O2=CO2+H2O and also for C7H16+O2 = CO2+H2O

OpenStudy (anonymous):

THANKS...ARE U TRYING ONLINE BALANCER? http://www.webqc.org/balance.php

OpenStudy (rane):

wht? no i am not trying anything i had to go outside

OpenStudy (rane):

then close your q'

OpenStudy (anonymous):

dont get upset please...i appreciate your kind help.....pl go ahead. answer

OpenStudy (rane):

5 C2H16 + 10 O2 = 10 CO2 + 4 H20 there u go

OpenStudy (rane):

did u get it?

OpenStudy (anonymous):

is it not C2H16+6O2=2CO2+8H2O

OpenStudy (rane):

i balanced it

OpenStudy (souvik):

is it \(C _{2}H _{16}\) ?

thomaster (thomaster):

Shouldn't it be \(\sf\large C_7H_{16}\) ?? \(\sf\large C_7H_{16} + 11 O_2 \to 7 CO_2 + 8 H_2O\)

OpenStudy (souvik):

that is what i am asking...!

OpenStudy (anonymous):

ya,it is C7H16...But, what is the actual chemical name?

OpenStudy (anonymous):

sorry Rane...it is c7h16

thomaster (thomaster):

\(\sf\large C_7H_{16}=heptane\)

OpenStudy (anonymous):

thanks thomaster

OpenStudy (anonymous):

can somebody enlighten me on 'adiabatic flame temperature'...may be some problems with solutions...